How many grams of silver will be generated if a 0.030A current is passed for 2.0 hours for the following reaction (not balanced)?
Ag+ (aq) + Cd (s) « Ag (s) + Cd2+ (aq)
Answer #1

Solukon! given chenicl Seaction g- Ca) 2 AgtCan) +es)2g+ given- time ( 2 hrs I = 0-03o Amp. / モ= 2メ60メ60gec. 72008e. A ccardThanks!